2. If you have a sample size of 20 cases, what is the minimum t-score needed to be statistically significant at the 0.01 level?
Answer #1

Solution:

We are given

n = 20

df = n - 1 = 20 - 1 = 19

α = 0.01

If test is lower tailed, then critical t value = -2.54.

If test is upper tailed, then critical t value = 2.54

Minimum t-score = 2.54

If test is two tailed, then critical t value = 2.86

Minimum t-score = 2.86